OpenStudy (anonymous):

find the volume of the solid obtained by rotating the region lying between the curves around the y-axis. y=sinx y=e^x x=pi/2 x=pi

OpenStudy (watchmath):

\(\int_{\pi/2}^{\pi }2\pi x(e^x-\sin x)\,dx\)

OpenStudy (anonymous):

can you help me with the steps?

OpenStudy (amistre64):

the graph doesnt makes sense ......

OpenStudy (amistre64):

you appear to have a boundary between boundaries

OpenStudy (amistre64):

or, that means that you spin it around the y axis; the part that is between ..got it

OpenStudy (amistre64):

if so; shelling is fine; 2pi {S} (xe^x - xsin(x)) dx ; [pi/2, pi]

OpenStudy (amistre64):

but it might be uglier than the disc method to integrate

OpenStudy (amistre64):

e^x =y x = ln(y); is ln^2(y) easy to integrate? havent tried it meself lol

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=int%28pi%28ln%28y%29%29^2%29dy+from+0+to+1 says its 2pi for that bit :)

OpenStudy (amistre64):

thnx :)

OpenStudy (anonymous):

i have couple more review questions can you help me amistre64?

OpenStudy (amistre64):

i can...maybe